OpenStudy (anonymous):

Evaluate the following limit: \[ \lim_{x \rightarrow -\infty} \frac{9x-4}{\sqrt{x^2+5x+3}}\] Current problem: please refer to my first comment here. Why should we add a negative sign for the denominator?

OpenStudy (experimentx):

try dividing both sides by x

OpenStudy (anonymous):

Near infinity the numerator is like 9x and the denominator is like x. How is the ratio look like?

OpenStudy (anonymous):

\[ \frac{9x}x=9 \]

OpenStudy (anonymous):

\[ \lim_{x \rightarrow -\infty} \frac{9x-4}{\sqrt{x^2+5x+3}}\]\[ =\lim_{x \rightarrow -\infty} \frac{\frac{9x-4}{x}}{\frac{\sqrt{x^2+5x+3}}{-\sqrt{x^2}}}\]Any problems so far?

OpenStudy (experimentx):

yep ... one way of looking it \[ \sqrt{x^2 + ..x + 222 .. } \approx x\] as x approaches infty

OpenStudy (experimentx):

Oops ... -inf ... didn't note. this hold for both case x-> -inf \[ \sqrt{x^2 + ..x + 222 .. } \approx x\]

OpenStudy (experimentx):

no probs ...so far.

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

I put a negative sign in front of \(\sqrt{x^2}\), why no problems for that? In other words, why should I do so?

OpenStudy (experimentx):

hold on ... now i'm confused. kinda forgot how to show it.

OpenStudy (experimentx):

if you make this approximation \( \sqrt{x^2 + ..x + 222 .. } \approx |x| \) then you will get answer

OpenStudy (anonymous):

I'm sorry but I haven't learnt that approximation, would you mind explaining it or..?

OpenStudy (experimentx):

let do this , change x = -a, and change the limits a->+inf

OpenStudy (anonymous):

I hope you don't mind me asking... But can we really do it in that way, changing the variable and the value it approaches?

OpenStudy (experimentx):

yep ... of course we can do ...

OpenStudy (experimentx):

this what we call substitution.

OpenStudy (anonymous):

Okay. So, a = -x \[ \lim_{x \rightarrow -\infty} \frac{9x-4}{\sqrt{x^2+5x+3}} = \lim_{a \rightarrow \infty} \frac{-9a-4}{\sqrt{(-a)^2+5(-a)+3}}= \lim_{a \rightarrow \infty} \frac{-9a-4}{\sqrt{a^2-5a+3}}\] Is that correct up to now?

OpenStudy (experimentx):

yep.

OpenStudy (anonymous):

\[= \lim_{a \rightarrow \infty} \frac{-9a-4}{\sqrt{(-a)^2-5a+3}}\]\[=- \lim_{a \rightarrow \infty} \frac{9a+4}{\sqrt{(-a)^2-5a+3}}\]\[=- \lim_{a \rightarrow \infty} \frac{\frac{9a+4}{a}}{\sqrt{\frac{(-a)^2-5a+3}{a^2}}}\]\[=- \lim_{a \rightarrow \infty} \frac{9+\frac{4}{a}}{\sqrt{1 - \frac{5a}{a^2}+\frac{3}{a^2}}}\]\[=\frac{-9}{\sqrt{1}}\]\[=-9\]Right?

OpenStudy (experimentx):

yep ...

OpenStudy (anonymous):

Hmm... Interesting... Would you mind explaining this step: change x = -a, and change the limits a->+inf ?

OpenStudy (anonymous):

I'm more interested in the part ''change the limits a-> +inf''. How does it work?

OpenStudy (experimentx):

let a = -x when x->0, a= ? when x->-inf, a= ?

OpenStudy (experimentx):

oh ... that's just clever manipulation i devised. the method is called ... anyhow get answer.

OpenStudy (anonymous):

Ahhh, that's clear to me now. Thanks! Should I change the limit to do similar question too?

OpenStudy (experimentx):

depending on the question ... and answer. just observe carefuly

OpenStudy (anonymous):

And also, what is the usual practice of doing such question? I saw the professor did in the way as I've typed in the first comment, but I'm not quite sure why she added a negative sign there. Her explanation was ''it's because when x-> -inf.''

OpenStudy (experimentx):

kinda true ... but changing variable is more cleverer.

OpenStudy (experimentx):

there a multiple method to approach problems.

OpenStudy (anonymous):

I'm not sure if changing variable is *accepted* (certainly it works this case as the answer matches), and I don't quite understand the professor's way to solve it. I would appreciate if you or anyone can help me understand. This has been a trouble to me for more than a year actually.

OpenStudy (experimentx):

look at the graph of both numerator and denominator ...

OpenStudy (anonymous):

graph?

OpenStudy (experimentx):

the numerator goes to -9 times infinity the denominator goes to + infinity --------------------------------- what would the ratio would be like?

OpenStudy (anonymous):

Can we really cancel the ''infinity'' ?

OpenStudy (experimentx):

hm .. not really. this doesn't go to infinity ... it rather tends to infinity

OpenStudy (experimentx):

that's how you get how a function behaves around some point. i like the concept of limit ... the whole calculus is based on it.

OpenStudy (anonymous):

tends to inf., but can we know what is near to infinity? (well, that's probably what we're looking for in limit though..)

OpenStudy (experimentx):

yep.

OpenStudy (experimentx):

1/x -> 0 as x-> inf 1/x behaves as 0 as x tries to go higher and higher.

OpenStudy (anonymous):

I got that concept. Just not sure of the negative sign. It's such a trouble.

OpenStudy (experimentx):

well ... try one of those two approaches ... also try discussing with your prof.

OpenStudy (anonymous):

Alright, thanks for your help. I'll just leave the question open to see if anyone can help. Thanks again for your time!

OpenStudy (experimentx):

yw